# Cardinality of Real Numbers and Set Axiomatization

### v2

@@ 3 Conclusion
Discussed are differences between various models that generate sets of numbers and how the inclusion of specific axioms influences the cardinality of these generated sets. Considered are some open questions and discrepancies among the current leading models, particularly those represented by V - ultimate L and those based on Martin’s axioms. Additionally, two explicit examples of generating real numbers are considered. It is demonstrated that expanding the $ZF$ model with the $AC$ highlights uncertainties in determining the cardinality of infinite sets. The provided examples are consistent with the Löwenheim-Skolem theorem. It is shown that the AC facilitates the denumerability of real numbers, and issues are identified within proofs of non-denumerability. The second example is also the first sequence representing all real numbers and the first explicit example of Wang’s constructive model.

### v3

@@ 6 Conclusion
Discussed are differences between various models for generating sets of numbers and how the inclusion of specific axioms influences the cardinality of these generated sets. Considered are some open questions and discrepancies among the current leading models, particularly those represented by V - ultimate L and models based on Martin’s axioms. Additionally, two explicit constructions of generated sets of real numbers are analyzed in detail.
It is demonstrated that expanding the $ZF$ framework with the Axiom of Choice exposes inherent uncertainties in determining the cardinality of infinite sets. The constructed sets, which are both everywhere dense and Dedekind complete yet countable, align with the Löwenheim-Skolem theorem. The role of $AC$ is emphasized in facilitating the denumerability of these generated complete sets, while tensions are identified between the Nested Interval Property and classical proofs of non-denumerability.
The second generated set serves as both a sequence representing all real numbers and the first explicit example of Wang’s constructive model. This dual role highlights the interplay between classical and constructive approaches to generating the continuum, offering a bridge between theoretical frameworks and concrete constructions.
These findings underscore the intricate relationship between axiomatic extensions, foundational properties such as completeness and density, and the nature of infinite sets. They provide a foundation for further exploration of the implications of specific axioms on set cardinality and the continuum hypothesis in modern set theory.